I don't understand on how to solve complex fractions with unlike denominators.
Please help.
4/3xyz^2+2/xyz-2/3z^2
Thank-YOu

Answer by hailu(13)

4/3xyz²+2/xyz-2/3z²
-multiply both the numerators & denominators with suitable terms, to get like denominators
=(4*1)/(3xyz²*1)+(2*3z)/(xyz*3z)-(2*xy)/(3z²*xy)
= (4/3xyz²)+(6z/3xyz²)-(2xy/3xyz²)
= (4+6z-2xy)/ 3xyz²
Unlike terms, so u don’t add them, which means, dats ur answer!
u r welcome